The pulse rates of adult men approach a normal distribution with a mean of 80 bpm(beats per minute) with standard deviation of 7 bpm. If 60 bpm to 100 bpm is known to be normal. How many percent of the adults have above or below normal pulse rates?


1
Expert's answer
2022-03-19T02:37:27-0400

Solution

Population mean "\mu=80"

Standard deviation "\sigma =7"


Percentage above "100" and below "60"

"Z=\dfrac{X-\mu}{\sigma}"


"Z_{60}=\dfrac{60-80}{7}=-2.86"


"Z_{100}=\dfrac{100-80}7=2.86"


From normal distribution tables

"P(below~ 60) =0.00212"


"P(above~100)=1-0.99788"

"P(above~100)=0.00212"


Total "=0.00424"


Percentage

"=0.00424\times100"

"=0.424\%"